The displacement of a particle is given by r(vector)=A(i^ coswt+j^ sinwt). The motion of the particle is:


a) an SHM


B) ON A straight line


c) on a circle


d) WITH CONTANT ACCELERATION

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Ramsing

r=A(i coswt+j sinwt)

x=Acoswt

y =Asinwt

 

square and add

x2 +y2 = (Acoswt)2 +(Asinwt)2

x2 +y2 = (A)2

this the equation of circle

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin